### v2

# Electrical spin manipulation in double SrTiO3/LaAlO3 quantum dots

Abstract
The spin dynamics in two electron double quantum dots embedded in two dimensional electron gas at the interface between SrTiO3 and LaAlO3 is studied by an exact numerical solution of the time-dependent Schrödinger equation in the context of the electric dipole spin resonance experiment. Based on the three band model of -electrons localized at Ti ions on the square lattice we analyze in details the singlet-triplet transition induced by the AC electric field, in the magnetic field range close to the avoided crossing which appears as a result of the spin-orbit coupling. For symmetric double quantum dots the single photon spin-flip transition is prohibited due to the parity symmetry and the transition can occur only by the higher order two-photon processes. For a weakly asymmetric system, when the first order singlet-triplet transitions are released due to the parity symmetry breaking, the spin-flip transition has a character of the Rabi oscillations for a low electric field amplitude. As the amplitude is increased the frequency of the transition is blueshifted (redshifted) for the magnetic field below (above) the single-triplet avoided crossing. For a sufficiently high magnetic field and high AC field amplitude the electric field drives the system across the avoided crossing inducing the spin-flip by the Landau-Zener-Stueckelberg-Majorana effect. Finally, the optimization of the geometrical parameters of the system with respect to the time of spin-flip of its fidelity is also presented.
I introduction
The control the spin of electrons confined in quantum dots (QDs) has attracted a growing interest in recent years, as it lies at the heart of developments towards a scalable spin-based quantum computer Burkard et al. (2023); Nadj-Perge et al. (2010). The efficiency of spin manipulation is entirely contingent upon the properties of the material platform, which dictate the strength of the spin-orbit coupling (SOC) Giglberger et al. (2007); Rashba and Efros (2003); Bychkov and Rashba (1984) or the hyperfine interaction with the nuclear spin bath. The latter is a source of spin decoherence Khaetskii et al. (2002) limiting the usability of a specific material platform in accordance with one of the Loss di Vincenzo criteria Loss and DiVincenzo (1998), whereby the quantum system must remain coherent for times much longer than the duration of elementary logic gates. For this reason, silicon-based QDs have been intensively studied over the last years due to their prolonged spin coherence time Mills et al. (2022).
With respect to that, another promising but still unexplored platform for QD-based spin qubits technology are transition metal oxide heterostructures, such as SrTiO3/LaAlO3 (LAO/STO) interfaces Joshua et al. (2012); Maniv et al. (2015); Monteiro et al. (2019); Pavlenko et al. (2012); Biscaras et al. (2012). The electronic properties of the two-dimensional electron gas (2DEG) formed at the LAO/STO interface is determined by the 3-orbitals Diez et al. (2015); Khalsa et al. (2013); van Heeringen et al. (2013); Wójcik et al. (2021); Zegrodnik and Wójcik (2020), implying that direct and indirect decoherence stemming from interaction with the nuclear bath can be significantly mitigated, as it is proportional to the square of the wave function at the nuclei position Khaetskii et al. (2002) - this value diminishes to zero for d-electrons. This property, combined with the strong spin-orbit interaction Caviglia et al. (2010); Yin et al. (2020), high mobility Ohtomo and Hwang (2004), and susceptibility to electrostatic gates at a level comparable to semiconducting materials Bjørlig et al. (2020); Thierschmann et al. (2018); Guenevere et al. (2017), initiates the currently ongoing research on LAO/STO-based quantum dots as potential spin-qubits with inherent scalability of 2D systems. The first experimental realization of electrostatically defined LAO/STO QDs has been already reported, yielding a Coulomb blockade diamond characteristic for a well-defined confinement Bjørlig et al. (2020); Thierschmann et al. (2018). The lateral confinement within LAO/STO interface was created earlier with the litography techniques Cen et al. (2009) and a detection of the confined charge island, or a quantum dot in a system tailored with nanolitography has been reported Maniv et al. (2016).
The demonstration of electrical spin manipulation in these systems constitutes the next milestone in spin-qubit realization based on the transition metal oxide heterostructures. Significant progress in this field has been recently achieved in electrostatically defined semiconducting QDs Schroer et al. (2011), where the microwave magnetic field, previously used to control spin dynamics Koppens et al. (2006, 2008), has been replaced by the AC gate voltage Nadj-Perge et al. (2010, 2010, 2012); Golovach et al. (2006). In this technique, called electric dipole spin resonance (EDSR), the microwave potential applied to one of the gates generates a time-dependent potential which, due to the momentum-dependent spin-orbit interaction Giglberger et al. (2007); Rashba and Efros (2003); Bychkov and Rashba (1984) and the resulting effective magnetic field drives transitions between hybrydized spin-up and spin-down states, leading to Rabi oscillations Koppens et al. (2006). For two electrons, the SOC hybridization induces level repulsion between the spin singlet and triplet states, which is often used to determine the strength and direction of the effective SOC field Stepanenko et al. (2012). Recently, EDSR technique applied to double quantum dots (DQDs) allowed for a demonstration of Pauli spin blockade Nowack et al. (2007); Nadj-Perge et al. (2010). In this regime, the system initially set to the spin triplet state , with tunneling from the left to the right QDs prohibited by selection rules, transmits to the spin singlet state , which is possible due to the SOC generated by the AC field applied to one of the gates Nowak et al. (2012).
Although spin manipulation in LAO/STO QDs has not yet been demonstrated experimentally, theoretical studies allow us to gain a first insight into the spin dynamics in these multiorbital systems. In our recent paper Szafran et al. (2024), we have discussed in detail the manipulation of electron spin in a single QD in the context of EDSR experiment. We have shown that for a single electron, the spin-flip in the ground state has the character of a Rabi resonance, while for two electrons, the singlet-triplet transition is forbidden by parity symmetry. The latter one is possible via a second-order, two-photon process, which exhibits a two-state Rabi character for low AC field amplitude.
Here we extend our study to the case of two-electrons in double quantum dots. Based on the time dependent scheme we simulate the singlet-triplet transition induced by the AC electric field. We consider the magnetic field range in the nearest of the avoided crossing which appears as result of the SO coupling. The calculated transition between singlet-triplet states, its characteristics, i.e. duration time and fidelity, are discussed with respect to the system asymmetry which is needed to induce the single photon transition by the parity symmetry breaking, as well as the coupling strength between the dots. Finally, the full symmetric system is also analyzed with respect to the higher order singlet-triplet transitions.
The present paper deals only with the electronic part of the system without the phonon-field. The coupling of electrons with the charge environment via acoustic phonons, that leads to the effects of relaxation and decoherence Fischer et al. (2009); Climente et al. (2007), needs to be integrated in a future model to account for these effects. The phonon field at the LAO/STO interface is a subject of only recent studies, in particular the effects of the surface on the optical phonon branch near the energy of meV has been determined in Ref. Liu et al. (2023).
The manuscript is organized as follows: in Sec. II we present a theoretical model of double quantum dots in 2DEG at the (001) LAO/STO interface as well as the scheme used for time-dependent two electron calculations. Sec. III contains the analysis of electronic spectrum of two-electron double quantum dots as well as results of time dependent simulations (EDSR), finally summary and conclusions are included in Sec. IV.
II Theoretical model
II.1 The single-electron problem
We consider 2DEG in the quantum well created at the (001) LAO/STO interface Popović et al. (2008); Pavlenko et al. (2012). The system is characterized by a significant atomic SO coupling and a strong Rashba interaction due to the asymmetry of the vertical electric field at the interface Diez et al. (2015); van Heeringen et al. (2013); Khalsa et al. (2013). We use the real space representation Szafran et al. (2024) of the tight-binding Hamiltonian Diez et al. (2015) spanned by orbitals of Ti ions that are arranged in a square lattice. The Hamiltonian of the system is given by
where is the vector of electron anihilation operators corresponding to states with the spin on the orbital at the position . In Eq. (II.1), accounts for the splitting of the orbitals degeneracy and the in-plane external potential defining the quantum dot confinement
where the tight-binding parameters meV, meV, meV, are adopted after Ref. Zegrodnik and Wójcik (2020); Diez et al. (2015), and is the external potential given at the ion position (.)
The atomic spin-orbit coupling is defined by and is given by
where the matrix corresponds to with the orbital angular momentum represented in the -orbitals basis and we take the SO coupling parameter meV Caviglia et al. (2010).
Finally, the coupling of the external magnetic field to the spin and atomic orbital momentum of electrons is taken into account by the Hamiltonian
where is the the Landé factor Ruhman et al. (2014); Bjørlig et al. (2020), is the Bohr magneton, with and with
The hopping elements in Eq. (II.1) consist of the kinetic term (spin-conserving intersite hopping) and the Rashba SOC (mixing between the and the two other orbitals) and take the form
and
with the Rashba SO coupling parameter meV Yin et al. (2020). Finally, the hybridization between , orbitals are taken into account by
with the hybridization parameter meV Diez et al. (2015).
The external magnetic field induces the Aharonov-Bohm phase shifts for the electrons moving along the interface, that we account for using the Peierls approach, i.e. multiplying the elements of the Hamiltonian corresponding to the hopping between ions placed at points and by a factor , where is the vector potential taken in the symmetric form for . Most of the results provided below are obtained for this perpendicular magnetic field orientation. However, the results for in-plane orientation are also provided. The in-plane field does not induce any Aharonov-Bohm phase shifts, produces no orbital effects and enters to the Hamiltonian only via the spin Zeeman term.
The in-plane external potential that models the double quantum dot system is assumed as a superposition of two Gaussian functions
where is the distance between the minima of the separate functions (centers of the left and right QDs), and define the depth of each potential minimum and nm is the size parameter of the separate Gaussians.
In our model we assume that the potential of each of the dots separately is rotationally symmetric. For the parameters applied below the tunnel coupling between the dots is non-negligible, so that the single-electron wave functions are extended over both the dots. Any deformation of the Gaussian potential with one of the dots will predominantly act as an effective potential shift in one of the dots. The shift would induce an asymmetry in the double dot potential that, as we show below, has a very pronounced effect on the dipole matrix elements.
II.2 The two-electron problem
The two-electron problem is defined by the Hamiltonian
with the dielectric constant Gariglio et al. (2015). The dielectric constant of STO can reach the values as large as 10 000 but can be limited in an electric field. As shown in the Appendix to Ref. Szafran et al. (2024) the electric field near the interface hosting 2DEG is very strong and leads to a reduction of the value to the one we apply in the present paper.
The problem is solved with the exact diagonalization method with the basis spanned by the atomic orbitals
where the summation runs over the position of ions , orbitals on the ion and the -component of the spin indexed by while is the spin-up or spin-down eigenstate. In the sums is equivalent to the triple of indeces and is one of the orbitals localized on the ion position . Evaluation of the Coulomb integrals has been performed in real space using the two-center approximation Guerrero-Becerra and Rontani (2014) and integration over the orbitals after Ref. Szafran et al. (2024). In the calculations, we use up to 50 lowest-energy single-electron states that produce 1225 Slater determinants as a basis for the two-electron problem.
II.3 Time stepping
We study the spin flips induced as singlet-triplet transitions between the eigenstates of the two-electron Hamiltonian with an external periodic perturbation of the potential . The time evolution is determined by integration of the time-dependent Hamiltonian . The solution is obtained in the basis of time-independent Hamiltonian eigenstates, with . Substituting this form of the wave function to the Schrödinger equation and projecting the resulting equation on the eigenstates one obtains a system of equations for ,
that we solved using the Crank-Nicolson time stepping with the two-electron ground-state as the initial condition, i.e. .
The present model does not take into account the relaxation and dephasing mechanisms. In a paper by two of the present authors Nowak and Szafran (2014) we considered the effects of the relaxation via acoustic phonons for EDSR experiments involving two states and demonstrated that the effect may lead to an off-resonant transition to a third state with a lower energy than the two states coupled through a resonant AC pulse. In Ref. Szafran et al. (2024) we argued that, when the considered transitions involve the two-lowest energy states in the spectrum, the relaxation mechanism is of a secondary importance. The dephasing in III-V and Si quantum dots is mainly due to the hyperfine interactions with the nuclear spins which should be absent in STO materials since the electron states are spanned by the 3d orbitals. The dephasing due to the interaction with the charge disorder in the surrounding matrix should be of a limited importance given the large value of the dielectric constant.
III Results
III.1 The two-electron eigenstates
In Fig. 1 we plotted the charge density of the lowest-energy spin singlet (a,c) and the lowest-energy spin triplet (b,d) for a symmetric quantum dots with spacing between the Gaussian minima of nm (a,b) and nm (c,d). The results for a stronger coupling [Fig. 1(a,b)] and in particular the form of the densities in the barrier region indicate that the singlet (triplet) charge density has a bonding (anti-bonding) character. For weaker interdot coupling [Fig. 1(c,d)] the charge densities in both spin states are more similar and the bonding (anti-bonding) character of the states is less pronounced.
The two-electron energy spectra for the considered DQDs are displayed in Fig. 2(a,b) for nm and in Fig. 2(c,d) for nm. At the ground-state is spin-singlet for any interdot distance . The singlet-triplet energy splitting (the exchange energy) at varies strongly with , e.g. the interdot tunnel coupling, and for nm it is equal to 185.5 eV, decreasing to only 33.1 eV for nm. The applied external magnetic field promotes the spin-polarized triplet for the ground-state. The singlet-triplet ground-state transformation occurs via an avoided crossing [Fig. 2(b,d)] that is opened by the spin-orbit coupling. The position of the avoided crossing at the magnetic field scale is determined by the exchange energy at . For nm ( nm) the center of the avoided crossing occurs at 1.05 T (0.19 T) with the width (e.g. minimal energy spacing between the anticrossing energy levels) equal to 12.3 eV for the stronger coupling and only 2.68 eV for the weaker coupling.
To account for an intrinsic or intentional asymmetry in the confinement potential of the double quantum dot system we introduce unequal depths of the Gaussian quantum wells forming the artificial molecule. In Fig. 3(a) we plot the exchange energy at as a function of the right quantum dot potential calculated for the left potential set at meV and nm (black line) as well as nm (red line). The exchange energy is minimal for the symmetric system. For an asymmetric double dot in the spin-singlet state both the electrons tend to occupy the deeper quantum dot [Fig. 3(b,c)] which lowers the singlet energy with respect to the spin triplet for which the double occupancy of the lowest orbital is forbidden by the Pauli exclusion principle. In contrast, the reaction of the charge localization to the asymmetry in the triplet state is weak [Fig. 3(b,c)]. Note, moreover, that a small asymmetry does not introduce a significant charge redistribution to the charge in the spin-singlet state which is a result of the electron-electron interaction. Only for nm or nm the electrons in the singlet start to occupy the deeper dot. For the weaker interdot coupling the transition to the deeper dot is more abrupt [cf. Fig. 3(b) and (c)].
For a symmetric potential the single-electron Hamiltonian eigenstates are also eigenstates of the parity operator , where is the scalar parity operator . The lowest-energy singlet and the lowest spin-polarized triplet that participate in the avoided crossing have the same – negative – parity – see Fig. 4. The same symmetry of the lowest singlet and triplet states has a direct consequence for the spin transitions driven by AC electric field, namely the dipole transition matrix elements between the two lowest-energy levels vanish so the first-order (single-photon) transitions is forbidden and only the second-order processes (two-photon) can occur. Similar behavior has been observed for two electrons in single quantum dot and discussed in details in our previous paper Szafran et al. (2024).
III.2 EDSR in weakly asymmetric system
Let us first study the case of weakly asymmetric system with meV. For small asymmetry the electron charge is still more or less evenly distributed over the dots, but the asymmetry lifts the parity selection rule that forbids the single-photon, first-order transitions.
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 2.58 | 2.54 | 0.45 | 0.4 |
2 | 0.041 | 0.23 | 0.31 | 0.15 |
3 | ||||
4 | ||||
5 | ||||
6 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 0 | 0 | 0 | 0 |
2 | 0.138 | 0.051 | 0.051 | 0.138 |
3 | 0.213 | 0.212 | 0.261 | 0.346 |
4 | 0.292 | 0.374 | 0.182 | 0.689 |
5 | 1.318 | 1.317 | 1.362 | 1.449 |
6 | 2.640 | 2.639 | 2.684 | 2.772 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | -1.94 | -1.98 | ||
2 | -1.99 | -1.96 | ||
3 | ||||
4 | 1.985 | 1.985 | 1.985 | 1.985 |
5 | ||||
6 |
Fig. 5(a) shows the spectrum for nm near the singlet-triplet avoided crossing that is shifted to higher magnetic field by the asymmetry due to the increased exchange energy [Fig. 3(a)] (to T from T in the case). The width of the avoided crossing is also enlarged: from 12.3eV (for meV) to 13.8 eV (for meV). The essential quantity for the first-order spin transitions is the dipole matrix element between the two-lowest energy levels which is plotted by the black curve in Fig. 5(b) and listed in Table I for selected values of the magnetic field. In Table I the first row corresponds to the average value of for the ground-state which are about 2.5 nm due to the shift of the wave function to the deeper dot. The shift gets smaller after the avoided crossing when the spin-polarized triplet replaces the spin-singlet in the ground-state. By Pauli exclusion principle the electrons in the triplet state cannot contain an admixture of configuration with both electrons in the single-electron ground-state of the deeper dot, hence the smaller shift to the deeper quantum dot for the triplet state. For an ideally symmetric system a similar singlet-triplet avoided crossing of energy levels is obtained with the interchange of the average values of the spin between the two lowest-energy levels, but the transition dipole matrix element is zero. For the system with lifted parity symmetry, the matrix element reaches its maximum in the center of the avoided crossing which allows the first order transition possible.
We studied the transitions from the ground-state to the first-excited state near the single-triplet avoided crossing induced by the external AC sine electric field, with the initial state set to the ground-state. The maximal occupancy of the first-excited state for the simulation lasting ns is given in Fig. 6 for T (a), T (b), T (c), and T (d) with the amplitude of the AC field increasing in steps of 5 V/nm.
Fig. 6 shows the maximal probability of the singlet-triplet transitions that are found during 5 ns simulation for a range of AC amplitudes and four selected perpendicular magnetic fields 0.434 T, 0.94 T, 1.5 T and 2 T. The matrix dipole elements for the singlet-triplet transitions, the energy above the ground-state and the z-component of the total spin for the four selected magnetic field values are given in Table I, II and III, respectively. The values are selected symmetrically with respect to the singlet-triplet avoided crossing – see second row of Table II. For the magnetic fields in Fig. 6(a,b) and Fig. 6(c,d) the ground-state is singlet and triplet, respectively. At the amplitude of 5 eV the largest peak corresponds to the singlet-triplet transition that occurs at the driving energy near the singlet-triplet energy difference. For T and T, where the dipole matrix element is large, at the lowest-amplitude, we also resolve the peaks of higher-order transitions at half and rd of the frequency for the single-photon transition processes. In an ideal symmetric system that we considered previously Szafran et al. (2024) the first-order singlet-triplet transitions are forbidden by the parity symmetry reasons. The transition matrix element vanish for the states that have the same parity as in the case of the singlet and the spin-polarized triplet in an inversion-symmetric system (cf. Fig. 4). The singlet-triplet transitions can still be induced for a symmetric system but with a subharmonic, two-photon transitions. The two-photon transitions at low AC frequency appear at the energy equal to roughly half the energy difference between the singlet and triplet states. The multiphoton transitions are higher order processes that require a much longer spin-flip time than the single-photon transitions. As the AC field amplitude is increased the frequency of the spin-flip transition [Fig. 6(a,b)] for the magnetic field below the single-triplet avoided crossing is blueshifted at higher AC field amplitudes. On the other hand the frequency for the spin transition for the magnetic field above the single-triplet avoided crossing gets redshifted for high AC amplitudes [Fig. 6(c,d)]. In both cases, in an intense AC field, we find an effective upbeat of the triplet energy with respect to the singlet.
Fig. 7 shows the time-resolved occupation of the ground-state singlet (black line) and the excited triplet (red line) as well as the overal occupation of higher excited states (blue) line for T for the amplitude of the AC field increasing from 10eV/nm (the highest row of plots) to 40eV/nm (the lowest row of plots). The evolution shows the results corresponding to Fig. 6(a), only longer evolution times than 5ns were taken when necessary. The left column of the plots in Fig. 7 shows the time evolution for the resonant single-photon frequency and the right column of the plots to the first subharmonic (two-photon) transition. The AC frequencies corresponds to the rightmost and second to the right peaks in Fig. 6(a). Note the difference of the times necessary to flip the spins in the single- and two-photon-transitions (left vs right column of the plots in Fig. 7). The blue line in Fig. 7 corresponds to the overall contributions of the higher energy-states. This contribution is nearly missing for the lowest amplitude considered in Fig. 7(a,b), and grows for higher AC amplitudes. Note, that the triplet occupation (the red line in Fig. 7) preserves its sine form also for higher AC amplitudes. On the other hand there is a fast oscillation that is superposed on the cosine form of the occupation line for the ground-state singlet. The ground-state singlet is strongly dipole-coupled to the 5th and 6th singlet states (see Table I) that are much higher in energy (Table II) with the large energy difference leading to a fast rate of these off-resonant transitions between the ground-state and excited singlets. The contribution of the higher excited states vanishes when the occupation of the ground-state singlet is zero. Figure 8 shows the overall contribution for all the excited states (black line) in the time evolution of Fig. 7(g) and the separate contributions from the 5-th and 6-th states (black line and blue line, respectively). The contribution of the 5-th state is larger consistently with the values of the dipole matrix element of Table I (see the first column of Table I corresponding to the perpendicular magnetic field of 0.434 T). Contribution of higher energy states is negligible.
The spectrum for the magnetic field increased from 0.434T to 0.94T [Fig. 6(b)] has much better resolved subharmonic structure due to times larger dipole matrix element (the second row of Table I and Fig. 5(b)). In Fig. 6(c) for 1.5T we find that the regular transition spectrum gets deteriorated at higher frequency with the disappearance of the single-photon peak. A near structure of peaks is restored at 2T – see Fig. 6(d). Both Fig. 6(c,d) in contrast to Fig. 6(a,b) develop a large values of spin-flip probability at the lowest range of the frequency. This is due to the electric field sweeping across the singlet-triplet avoided crossing. In order to explain this effect in Fig. 9 we show the two lowest-energy states for the double dot system with a time-independent electric field of added to the potential. The avoided crossing for T [Fig. 9(b)] is reached at V. Note, that in AC field with the amplitude of V/nm the maximal occupancy of the singlet-state is no longer zero at low limit [cf. Fig. 6(c)]. At higher amplitude the low-frequency limit in Fig. 6(c) corresponds to an adiabatic sweep of the potential across the singlet-triplet avoided crossing with 100% spin flip probability in a limit of an adiabatic sweep.
For T and T at higher amplitude the AC field drives the system across the singlet-triplet avoided crossing so that the process for the spin-flip can acquire the Landau-Zener-Stueckelberg-Majorana (LZSM) character Ivakhnenko et al. (2023). The ground-state system that is driven across the avoided crossing can either stay in the ground-state (adiabatic limit) with the transformation of the wave function from the one characteristic to the ground-state at the other side of the avoided crossing or preserve its wave function and find itself in the excited state. For periodic modulation the phases for both processes accumulate or cancel that leads to the LZSM interference pattern Ivakhnenko et al. (2023). The interference produces the comb visible in Fig. 6(c) at higher amplitude and low frequency. For T we reach a similar point at V/nm [cf. Fig. 6(d) and Fig. 9(c)]. The change of the character of the triplet-singlet transition is better illustrated by the transfer probability plotted as a inverse of the driving energy presented in Fig. 10 for T. As the amplitude increases from 5V/nm to 25V/nm higher-order peaks corresponding to multiphoton transitions appear but the corresponding peaks fall below 100% at large . For V/nm and above, the transition probability peaks acquire the height of 100% that does not fall below at large , and appear periodically on the scale, which are characteristics of the LZSM interference Ivakhnenko et al. (2023).
The time-resolved occupation of the states involved in the system dynamics for the magnetic field of 2T is illustrated in Figs. 11 and 12. Fig. 11 corresponds to the one- and two-photon transitions. At the AC amplitude of V/nm [Fig. 11(a,b)] the occupation has a similar character to the ones found for T [cf. 7(a,b)]. At higher amplitude the contribution from the higher excited singlet states (blue line in Fig. 11(e-h)) is correlated with the red line – that now corresponds to the spin-singlet which is the first excited state in the field of 2T. For the amplitude of 20V/nm the system is still not driven across the avoided crossing, only approach it (see Fig. 9(c)). When this happens (cf. Fig. 11(e,f) and (g,h) for the amplitude of the AC field of 30 and 40V/nm, respectively) the time evolution of the occupations deviates visibly from the simplest periodicity – see the shapes of the maximal occupation probability at each turn. Note, that the time evolution in Fig. 7 for the magnetic field of 0.434 T, when the sweep across the avoided crossing does not occur, is periodic in the entire range of frequencies considered. In Fig. 11 for higher AC amplitudes the second order transition becomes nearly as fast as for the two-photon transition which never occurs for the driving off the avoided crossing. The driving in Fig. 11 corresponds to the first two peaks at right of Fig. 6(d), i.e. to a relatively large AC frequency. The time evolution for the lowest-frequency i.e., rightmost peak in Fig. 10 is given in Fig. 12(a). The orange line in this peak shows the curve that modulates the external field. The center of the AC of Fig. 9(c) is reached at . Above this value the occupation of the ground (black line) and first excited (red line) state in Fig. 12(a) is nearly constant. Variation of the occupation is found when the function acquires values lower than . In Fig. 12(a) transitions from the ground to the excited state after each sweep of the AC field accumulate which results in a constructive LZSM interference. For contrast, in Fig. 12(b) we took the driving frequency between the two rightmost peaks of Fig. 10. Here the effects of the transitions do not accumulate but cancel at each sweep.
Figure 13(a) shows the position of the first order transition on the energy scale as a function of the AC amplitude. Note that, the magnetic fields considered in Figs. 13(a) and Fig. 6 were chosen to produce equal singlet-triplet energy splitting before and after the avoided crossings, hence the coalescent lines at the limit of zero amplitude. The frequency shifts before and after the singlet-triplet avoided crossing have opposite signs but similar magnitudes. The line for T is interrupted above 35 V/nm, since the resonance line can be hardly recognized in the spectrum [cf. Fig. 6(c)].
A strong AC amplitude induces appearance of other states than the lowest-energy couple in the time evolution. The maximal share of the higher energy states as a function of the amplitude is given in Fig. 13(c) as the leakage of the wave function outside the lowest-energy singlet and triplet. The dependence was determined at the first-order resonant transition and as we can see the leakage on the amplitude is similar for all the magnetic fields considered. At the amplitude of V/nm the maximal share of the higher energy states is about 2.5%.
The dependence of the spin-flip time on the amplitude for the first-order transition is plotted in Fig. 13(b). The minimal flip time of about 245 ps for T is obtained at the amplitude of 25 V/nm. For even stronger amplitudes the flip time no longer decreases which is related to the potential sweeps across the singlet-triplet avoided crossing and the leakage of the wave function to higher energy states for which the evolution no longer have a two-state Rabi resonance character Khomitsky et al. (2020).
Due to the very strong spin-orbital interactions in the present system the minimal EDSR-induced slip flip time of 245 ps is very short. Note that the fastest single-triplet transition is obtained when the time evolution is not longer a two-level Rabi cycle with the presence of higher-excited states of about 5% (see Fig. 13(c)). However, even for the lowest considered AC field amplitude, when the spin transitions follow the Rabi cycle, the spin flip times of a few nanoseconds are still very low as compared to the values observed in quantum dots defined in GaAs or Si. The corresponding electron spin inversion time mediated via the spin-orbit coupling in electrostatic double quantum dots in GaAs is 106 ns Nowack et al. (2007). Faster spin-inversion times of about 50 ns have been achieved with the microwave chip generating magnetic field locally in the GaAs quantum dots Koppens et al. (2006) in an electron spin resonance setup. For GaAs quantum dots with the conduction band formed by s-type electron orbitals the spin inversion can also be achieved in the random field of nuclear spins Laird et al. (2009) but the Rabi oscillations in the hyperfine field are very slow with the characteristic time of about 1 microsecond. The effects of the hyperfine interactions should be negligible in the present system due to the d-type of the Ti orbitals that span the space for the confined states. In silicon the hyperfine field can be minimized in isotopically enriched 28Si and the spin-flip time mediated by spin-valley coupling is of the order of 0.5 microsecond Klemt et al. (2023). Advanced quantum registers of spin-qubits in silicon quantum dots are characterized by faster spin-flip times of ns Philips et al. (2022). These devices Philips et al. (2022) are equipped with integrated micromagnets that form a gradient of the magnetic field which mediates the spin flips in external AC electric field that sweeps the confined electrons through the inhomogenous magnetic field region.
III.3 Orientation of the fields and the interdot spacing
In the paper we assume that the magnetic field is oriented perpendicular to the STO/LAO interface and that the AC electric field is oriented parallel to the axis of the double dot system. For a choice of the orientation within the plane of confinement the orbital effects of the magnetic field are absent. The energy spectrum for the weakly asymmetric system of Fig. 5 and the magnetic field oriented parallel to the axis is displayed in Fig. 14(a). One notices that for the in-plane magnetic field orientation the curvature of the energy levels as a function of is missing outside the singlet-triplet avoided-crossing. The avoided crossing is shifted to higher values of the magnetic field by about 0.1T. The spin polarization of the lowest-states as well as the dipole matrix element given in Fig. 14(b) are very similar to the results in the perpendicular magnetic field (Fig. 5(b)) with the exception of a small shift of the maximum due to the shift on the scale of the avoided crossing. The values of the matrix elements, the energy splittings and the spin components are given in Tables IV, V and VI, respectively. Overall, the orientation of the magnetic field turns out to be of a secondary importance.
The AC field oriented in the direction chosen in this paper produces large dipole matrix elements due to the spatial extension of the charge system (see Fig. 1). The field oriented parallel to the axis produces the matrix elements which are by two or three orders of magnitude smaller (Table VII). The present 2D model does not allow for a straightforward incorporation of the AC electric field variation in the direction perpendicular to the plane of confinement. However, as we argued in Ref. Szafran et al. (2024) the EDSR can be understood as a translation of the electron motion induced by AC field to an effective magnetic field that is – as in Rashba type coupling – orthogonal to the direction of the induced electron spatial movement. The effective field needs to be orthogonal to the external magnetic field polarizing the electron spins. Therefore, the AC modulation of the electric potential perpendicular to the plane of confinement should lead to the transitions between the states polarized in this direction. We do not expect the effect to be stronger than for the field oriented along the axis of the dots due to the very strong confinement of the wave functions in the direction (see Fig. 14 of Ref. Szafran et al. (2024)).
The effect of the distance between the Gaussian quantum well centers on the dipole matrix element and the singlet-triplet avoided crossing is given in Fig. 15(a,b). The curves in Fig. 15 are denoted by an integer , where the spacing between the centers is taken as where nm is the lattice constant. The values of nm and 31.2 nm used in the precedent section correspond to and , respectively. The interdot spacing has a pronounced effect on the position of the singlet-triplet avoided crossing that produces the maximum of the dipole matrix element. For a large interdot spacing the singlet and triplet tend to degeneracy that occurs for strictly separated electrons. A small magnetic field promotes the spin-polarized triplet to the ground-state. For a small values of the systems tends to a single elongated quantum dot and a large magnetic field is required for a ground-state spin transition. The maximum of the dipole matrix element is a non-monotonic function of the interdot spacing. In the limit of decoupled quantum dots with absent interdot tunnel coupling the dipole matrix decreases. On the other hand for a small value of the dipole matrix element cannot reach large values due to a limited extension of the wave functions. The magnetic field that is chosen for the spin manipulation needs to be placed off the avoided crossing since the spin states are evenly mixed at the center of the avoided crossing. Note, that the width of the peaks in Fig. 15(a) grows with decreasing which will affect the experimental workpoint.
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 2.59 | 2.58 | 0.52 | 0.43 |
2 | 0.042 | 0.193 | 0.46 | 0.18 |
3 | ||||
4 | ||||
5 | ||||
6 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 0 | 0 | 0 | 0 |
2 | 0.142 | 0.0598 | 0.0363 | 0.116 |
3 | 0.213 | 0.214 | 0.0245 | 0.329 |
4 | 0.284 | 0.367 | 0.493 | 0.656 |
5 | 1.318 | 1.318 | 1.352 | 1.433 |
6 | 2.640 | 2.640 | 2.674 | 2.755 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | -1.905 | -1.981 | ||
2 | -1.989 | -1.970 | ||
3 | ||||
4 | 1.992 | 1.192 | 1.992 | 1.991 |
5 | ||||
6 |
0.434T | |
0.94T | |
1.5T | |
2T |
III.4 Optimization of the dipole matrix moment
In the Rabi oscillations regime the rate of the first-order spin-flip transition is proportional to the value of the dipole matrix element. The latter is strictly zero for an ideally symmetric double quantum dots. In the precedent section we described the results obtained in a system with a weak asymmetry but the value of the matrix element can be optimized with the system parameters.
The dipole matrix element for nm is displayed in Fig. 5(b). In a system with larger interdot barrier ( nm), but still with the small asymmetry meV as in Fig. 5(b), the singlet-triplet avoided crossing occurs at lower magnetic field of T [cf. Fig. 16(a)] than for the stronger interdot coupling due to a reduction of the exchange energy [Fig. 3(a)]. The width of the singlet-triplet avoided crossing is 3.2 eV for nm. The maximal value of the dipole matrix element in the center of the avoided crossing is about nm, e.g. about 3 times lower than for nm [Fig. 5(b)].
The position of the singlet-triplet avoided crossing and the maximal value of the dipole matrix element strongly depends on the asymmetry of the confinement potential. In Fig. 16(b) and (c) we plotted the average spins and the dipole matrix element near the singlet-triplet avoided-crossing for nm with meV (b) and meV (c). For meV [Fig. 16(b)] the maximal value of the dipole matrix element is 40 times larger than for the small asymmetry meV [Fig. 16(a)] and about 4 times larger than for the large asymmetry meV [Fig.16(c)].
The solid lines in Fig. 17(a) show the absolute value of the dipole element at the center of the singlet-triplet avoided crossing as a function of the right-dot potential . The asymmetry of the confinement potential pushes the avoided crossing to higher values of the magnetic field (dashed lines referred to the right axis). The dependence of the matrix element on the asymmetry is non-monotonic which is related to the charge distribution in the singlet and triplet states. Fig. 17(b) shows the charge in the left and right quantum dot calculated for T – near the value of the magnetic field for which the maximum of the dipole matrix element is achieved in Fig. 17(a) for meV. The gap in the lines of Fig. 17(b) corresponds to the region of the singlet-triplet avoided crossing as a function of that we skipped for the clarity of the plot. At the left (right) hand side of the avoided crossing the ground-state for T is the spin triplet (singlet).
Comparing Fig. 17(a) and Fig. 17(b) we can conclude that the maximal value of the dipole matrix element is achieved for which on the one hand localizes the singlet entirely in the deeper dot but for which the triplet charge distribution is still evenly distributed in both dots. In the region of the avoided crossing the two-lowest energy states have hybridized wave functions that due to the contribution of the triplet are delocalized over both dots which produces a large value of the dipole matrix element. For large , on the other hand, both singlet and triplet states are confined within a single deeper quantum dot with potential perturbed by the potential of the left Gaussian. Due to this perturbation the dipole matrix element is non-zero. Since the system is localized in a single dot the matrix element does not achieve values as large as for the delocalized triplet state. Comparing Fig. 3(b) and Fig. 17(a) we also notice that at higher the triplet state is more easily localized in the deeper dot: for meV the charge in the right dot is at B=0, but already at T.
For the studies of the transition times we chose meV – near the maximum of the dipole matrix element [Fig. 17(a)] and three values of the magnetic field after the singlet-triplet avoided crossing, where the ground-state is the spin triplet – T, T and T – see Fig. 16(b). The singlet-triplet transition times as functions of the AC amplitude is given in Fig. 18. The transition times reach minima for lower values of the AC field amplitude than in Fig. 13(b) and start to grow for the values that exceeds the one for the LZSM transition regime. The LZSM regime is open for the amplitudes of V, 2.7V and 9.5 V/nm for T, T and T, respectively. Note, that for each value of the magnetic field the minimum spin flip time of about 330 ps is achieved for some value of the AC field amplitude. Since the amplitudes of the AC field near the minimal spin flip time are smaller than in Fig. 13(b), so are the wave function leakage values near the minimum which are only 0.5%, 0.02%, 0.0043% for V/nm, V/nm, and V/nm, corresponding to T and T, and T, respectively.
III.5 Spin flips for an ideally symmetric system
For completeness we present the results for an ideally symmetric system, for which the direct spin-flip transition is forbidden by the selection rule. Fig. 19 shows the probability of the spin-flip-singlet for nm and the external magnetic field set at T, for a simulation lasting ns and starting from the ground-state triplet as the initial state. The subsequent lines at the upper part of Fig. 19 correspond to a growing amplitude of the AC field. The most pronounced peak that is observed in Fig. 19 at weak AC field amplitude occurs at roughly half the singlet-triplet energy splitting that at the field of T is equal to meV [cf. Fig. 2(b)]. The spin-flip occurs for AC frequency set at this peak by a two-photon second-order process. Also, no third-order (three-photon) peak at 1/3 of is observed. For higher amplitude the position of this peak gets red-shifted and the peaks corresponding to even denominators, e.g. 1/4 and 1/6 of are formed.
The study of the AC driven spin-flips in a symmetric system with nm is summarized in Fig. 20 for two values of the external magnetic field T and T, both corresponding to the ground-state triplet. The position of the resonant second-order transition on the energy scale is redshifted with the amplitude [Fig. 20(a)]. In the studied range of the AC amplitudes – the spin-flip time is a decreasing function of the amplitude [Fig. 20(b)]. For T the LZSM regime opens at 45 eV/nm, and for T at still higher amplitudes which are outside the range studies in Fig. 20.
IV Summary and conclusions
We have studied two electrons confined in a lateral double quantum dot defined within the two-dimensional electron gas on the (001)-oriented LAO/STO surface. Utilizing the three-band model of -electrons localized at Ti ions on a square lattice, we have analyzed the energy spectrum of the system, paying particular attention to the singlet-triplet avoided crossing induced by the SO interaction. The calculated eigenstates have been employed as initial conditions for time-dependent calculations. In the magnetic field range close to the avoided crossing, the spin-flip singlet-triplet transition induced by an AC electric field has been analyzed in details in the context of the electric dipole spin resonance experiment.
We have demonstrated that for a fully symmetric system the first order singlet-triplet transition are forbidden due to the parity symmetry of the Hamiltonian which results in zeroing of the dipole matrix elements. To induce the spin-flip transition we introduce the asymmetry in the confinement potential in the form of unequal depths of the Gaussian quantum wells. Our simulations for a weakly asymmetric system show that the spin-flip singlet-triplet transition has the character of the Rabi oscillations for a low AC field amplitude, with minimal flip time of about 245 ps. Interestingly, when we increase the AC filed amplitude the frequency of the transition is blueshifted (redshifted) for the magnetic field below (above) the single-triplet avoided crossing. Moreover, we have noticed that for a magnetic field above the anticrossing the regular transition spectrum is highly deteriorated at a low frequency regime and the high AC field amplitude. This behavior has been explained as resulting from the Landau-Zener-Stueckelberg-Majorana transitions between the ground and first excited states which occur in the system when the electric field is large enough to drive the system across the avoided crossing.
The optimization of system geometrical parameters shows that the appropriate potential asymmetry between the dots highly reduces the leakage of transition to the higher states which can reach the value of about 0.05% with the spin flip time at the order of ps. Finally our simulations for a fully symmetric system provide insight into the possibility of the spin-flip transition by the multiphoton processes. The singlet-triplet transition studied here will be relevant for future EDSR experiments aimed at reading the spin state of the dot and paving the way to a spin-orbit qubit at the oxides interface.
V ACKNOWLEDGEMENT
This work is financed by the Horizon Europe EIC Pathfinder under the grant IQARO number 101115190 titled "Spin-orbitronic quantum bits in reconfigurable 2D-oxides" and partly supported by program „Excellence initiative – research university” for the AGH University. Computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) within computational grant no. PLG/2024/017175 was used.